OpenStudy (anonymous):

If an object is launched straight up into the air from a starting height of h_{0} feet, then the height of the object after t seconds is approximately h=-16t^2+v_{0}t+h_{0} feet, where v_{0} is the initial velocity of the object. Find the starting height and initial velocity of an object that attains a maximum height of 412 feet five seconds after being launched.

OpenStudy (anonymous):

h=-16t^2+v_{0}t+h_{0} => final velocity = h'(t) h'(t) = -32t + v(0) Maximum height is achieved after 5 seconds: => h'(5) = -32*5 + v(0) => v(0) = h'(5) + 32*5 = h'(5) + 160 Remember maximum height => h'(t) = 0. i.e. final velocity must be zero at maximum height just before the object starts decending. => v(0) = 160 + 0 = 160 Initial velocity = 160 b) Starting height: h=-16t^2+v_{0}t+h_{0} we know that v(0) = 160 ( from above) => h=-16t^2+v_{0}t+h_{0} => h=-16*25+ 160*5 + h_{0} => h =-400+ 800 + h_{0} But h = 412 ( given) => 412 = 400 + h_{0} =>h_{0} = 12 Initial height =12 Initial velocity = 160

OpenStudy (anonymous):

FYI you can prove the above plugging h(0) and v(0) values into the original equation : h=-16t^2+v_{0}t+h_{0} This should give the value of h as 412 concurring with the value given in the question. The trick was: a) differentiate the original equation to give the final velocity. b) set that final velocity to zero to get the original velocity.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

vw